# EBC-ZIP: Improving Blockwise Crowd Counting with Zero-Inflated Poisson Regression

### v1

@@ Abstract
Density map estimation has become the mainstream paradigm in crowd counting. However, most existing methods overlook the extreme sparsity of ground-truth density maps. In real-world crowd scenes, the vast majority of spatial regions (often over 95%) contain no people, leading to heavily imbalanced count distributions. Ignoring this imbalance can bias models toward overestimating dense regions and underperforming in sparse areas. Furthermore, most loss functions used in density estimation are majorly based on MSE and implicitly assume Gaussian distributions, which are ill-suited for modeling discrete, non-negative count data. In this paper, we propose EBC-ZIP, a crowd counting framework that models the spatial distribution of counts using a Zero-Inflated Poisson (ZIP) regression formulation. Our approach replaces the traditional regression loss with the negative log-likelihood of the ZIP distribution, enabling better handling of zero-heavy distributions while preserving count accuracy. Built upon the recently proposed Enhanced Block Classification (EBC) framework, EBC-ZIP inherits EBC’s advantages in preserving the discreteness of targets and ensuring training stability, while further improving performance through a more principled probabilistic loss. We also evaluate EBC-ZIP with backbones of varying computational complexity to assess its scalability. Extensive experiments on four crowd counting benchmarks demonstrate that EBC-ZIP consistently outperforms EBC and achieves state-of-the-art results. Our code is available at: https://github.com/Yiming-M/EBC-ZIP
Index Terms:
crowd counting, zero-inflated Poisson regression, scalability, blockwise prediction, real-time crowd counting
I Introduction
@@
Crowd counting aims to automatically estimate the number of people present in an image or a video frame. In recent years, it has attracted increasing attention due to its broad range of real-world applications, including public safety management [1], urban planning [2], large-scale event monitoring [3], and intelligent transportation systems [4].
Figure 1: Distribution of blockwise count values (block size = 16) across four crowd counting datasets: ShanghaiTech A & B [5], UCF-QNRF [6], and NWPU-Crowd [7]. The vast majority of blocks have a count of 0, highlighting the extreme sparsity of ground-truth blockwise density maps in these datasets.
While directly regressing the total crowd count from an image inherently discards essential spatial information about where people are located, most modern crowd counting methods [5, 8, 9, 10, 11, 12, 13, 14, 15] address this limitation by predicting density maps instead. However, most of them ignore the extreme sparsity of ground-truth density maps. In real-world crowd scenes, the vast majority of spatial regions contain no people, resulting most values of the ground-truth density maps being zero. If this severe imbalance is not properly addressed, it leads to two significant issues. First, models trained with standard regression losses (e.g., MSE or MAE) become biased toward predicting low or zero values, reducing their sensitivity to sparse yet informative regions. Second, the dominance of zero-valued regions weakens the supervisory signal during training, impeding the model’s ability to learn fine-grained density variations.
To address this issue, the majority of the existing methods [5, 9, 10, 11, 16, 17, 13, 18, 15] adopt blockwise prediction, which aggregates counts over fixed-size patches. Although it can alleviate this issue to some extent, it introduces a trade-off: Small block sizes provide little benefit in sparsity reduction, and large block sizes result in the loss of spatial resolution. As illustrated in Fig. 1, even with a block size of 8, the majority of blocks in most crowd counting datasets still have zero count. There have also been some efforts [19, 20, 18] to address the sparsity by training a segmentation model which aims to separate regions containing people from the background. However, most crowd counting datasets do not provide segmentation masks, they have to rely on Gaussian-smoothed ground-truth density maps to generate pseudo ground-truth segmentation masks. This introduces noise into the pseudo segmentation masks, as the head sizes are not annotated in crowd counting. Guo et al.[18] exploited the box annotations provided by NWPU-Crowd [7] to handle this problem, but compared with point annotations, box annotations are much more expensive to acquire. Besides, these methods introduce an extra task to counting, and how to balance the loss terms in such a multi-task setting poses a new challenge.
On the other hand, most loss functions used by existing crowd counting models are majorly based on blockwise mean squared error (MSE) of density maps. These loss functions, however, are grounded in implicit assumptions that the blockwise count values follow Gaussian distributions. Such assumptions are suboptimal as blockwise count values are inherently discrete, non-negative, and often heavily skewed toward zero. A more suitable statistical model for count data is the Poisson distribution, which captures both the discreteness and non-negativity of counts. Nevertheless, standard Poisson regression struggles with the overwhelming presence of zero counts in crowd counting tasks.
Figure 2: Scalability of the proposed EBC-ZIP framework. The $x$-axis represents the number of model parameters (in millions), and the $y$-axis indicates the Mean Absolute Error (MAE; lower is better) on the validation set of the ShanghaiTech Part B dataset [5]. Each circle represents a specific model, where the radius corresponds to the number of floating point operations (FLOPs) required to process a single $1920\times 1080$ image during inference. EBC-ZIP models (in blue) exhibit a favorable trade-off between model size and accuracy, demonstrating better scalability compared to other lightweight (yellow) and heavyweight (green) models.
To address the above mentioned issues, we propose to adopt a Zero-Inflated Poisson regression (ZIP) model, which explicitly distinguishes between two types of zero counts:
  • Structural zeros, which are hypothesized to be regions that are deterministically zero due to their semantic irrelevance. These include typical background areas such as the sky and buildings, as well as non-head body parts and non-central head regions  [^1] In most crowd counting datasets, only head centers are annotated. As a result, non-head regions like torsos and limbs receive zero count and are treated as structurally zero..
  • Sampling zeros, which refer to head-center regions that nonetheless receive zero count due to the limitations of point-based annotation and annotator subjectivity.
Since there structural and sampling zeros are not annotated separately, our ZIP model learns to distinguish these two types of zeros implicitly by optimizing the negative log-likelihood of the ZIP distribution. Unlike segmentation-based methods [19, 20, 18], our ZIP-based method offers several key advantages. First, we do not require any segmentation masks, and thus, no noise is introduced due to head size uncertainty and we do not need extra work on box annotations. Second, by modeling count as a zero-inflated Poisson distribution over image blocks, our framework captures structural sparsity without assuming the presence of clear object boundaries, making it better suited for highly occluded and densely crowded scenes. We choose to implement our ZIP model within the Enhanced Block Classification (EBC) framework [15], as compared with direct blockwise regression, EBC offers better training stability and preserves the discrete nature of count data. The integration of ZIP further improves this framework by allowing it to capture both structural sparsity and count uncertainty. We also investigate the scalability of EBC-ZIP by testing its performance using backbones with different computational complexities, such as MobileNet [21], MobileCLIP [22], CLIP-ConvNeXt [23, 24]. Fig. 2 highlights the scalability of our EBC-ZIP framework, which outperforms existing models [8, 16, 25, 26, 11, 27, 13] in both accuracy (MAE) and computational cost (FLOPs) across different parameter scales.
In summary, the main contributions of our work are as follows:
  • We propose a novel crowd counting framework EBC-ZIP based on Zero-Inflated Poisson (ZIP) regression, which explicitly models the extreme sparsity of ground-truth count distributions by disentangling structural zeros from sampling uncertainty. Built on top of the EBC framework, EBC-ZIP significantly improves its performance by replacing heuristic classification with a principled probabilistic formulation.
  • We present the first systematic study of framework-level scalability in crowd counting. Our EBC-ZIP framework generalizes well across a wide range of model sizes and architectures, from lightweight convolutional neural networks (CNNs) to vision-language models, achieving state-of-the-art performance under varying computational constraints.
  • We conduct extensive experiments on four crowd counting benchmarks to validate the effectiveness of our EBC-ZIP framework. Results show that our base model, EBC-ZIP-Base, consistently outperforms existing state-of-the-art methods across four different datasets.
II Related Work
II-A Handling Sparsity
To address the extreme sparsity of ground-truth density maps, many existing methods [5, 9, 10, 16, 17, 13, 18] adopt a combination of Gaussian smoothing and blockwise prediction. These approaches typically preprocess the binary annotation map by convolving it with Gaussian kernels, either of fixed or adaptive size. Specifically, several works [9, 10, 17, 13, 18] use fixed kernel bandwidths, while others [5, 16] adjust the kernel size according to the distance to the nearest annotated head. In principle, the kernel size should reflect the actual head size in the image; however, such information is typically unavailable in most crowd counting datasets. As a result, Gaussian smoothing inevitably introduces noise into the ground-truth density maps, which can degrade the model’s generalization ability [11]. To overcome these limitations, recent works [11, 15] propose to eliminate Gaussian smoothing altogether and instead rely solely on blockwise prediction. Empirically, these smoothing-free blockwise methods have demonstrated better performances compared to those using Gaussian-smoothed density maps. Nevertheless, the sparsity issue persists: as shown in Fig. 1, even with a block size of 8, which is used in the aforementioned blockwise prediction methods, over 85% of the blocks still have zero count in most crowd counting datasets. This extreme imbalance presents significant challenges for training and optimization, as the vast majority of the model’s supervision signal comes from zeros.
II-B Modeling Count Distributions in Density Maps
Most density-based crowd counting methods typically frame the task as a regression problem, where the model is trained to minimize the blockwise mean squared error (MSE) between the predicted and ground-truth blockwise density maps [5, 9, 28, 10, 16, 13, 18]. This loss function implicitly assumes that each blockwise count follows a Gaussian distribution centered at the predicted value. Although this formulation is simple and effective in many scenarios, it fails to capture the fundamental characteristics of count data, which is discrete, non-negative, and often exhibits a long-tailed distribution. DMCount [11] addresses a different issue by reformulating density estimation as an Earth Mover’s Distance (EMD) problem. By minimizing the Wasserstein distance between predicted and ground-truth density maps, DMCount explicitly models the global transportation cost needed to match the predicted mass to the annotated points. This approach improves robustness to spatial misalignment and annotation noise. However, similar to MSE-based methods, it still operates entirely in a continuous density space, treating the density maps as divisible mass rather than integer-valued counts. Also, the exact computation of EMD scales poorly with the number of bins ($\mathcal{O}(n^{3})$). In contrast, Bayesian Poisson Regression (BPR) is [29] an early attempt to model count from a discrete perspective. In BPR, an image is segmented into large regions, each assumed to follow a Poisson distribution, with the log-mean regressed from hand-crafted low-level features. While this formulation respects the discrete and non-negative nature of count data, the spatial segmentation is often coarse, and as a result, the independence assumption between regions is unrealistic in densely crowded scenes, where spatial correlation is strong. This issue can be mitigated by adopting a finer-grained modeling strategy by treating small blocks (e.g., $8\times 8$ or $16\times 16$) as the basic counting units. Local interactions tend to be weaker under this circumstance, so the assumption of blockwise independence becomes more plausible.
In this work, we propose to model blockwise count distributions using zero-inflated Poisson (ZIP) regression. Unlike existing methods, the zero-inflated Poisson (ZIP) distribution better aligns with the empirical distribution of blockwise crowd counts, as it naturally captures their discreteness, non-negativity, long-tailedness, and zero-heaviness.
III Method
Figure 3: Overview of the proposed EBC-ZIP framework. Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a backbone network (e.g., MobileNet [21], MobileCLIP [22], or CLIP [30]) extracts feature maps, which are shared by two parallel branches. In the EBC Poisson branch (top), the features are passed to a $\boldsymbol{\lambda}$-head with a softmax activation, producing the blockwise count distribution $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ over $n-1$ positive bins. The Poisson rate per block is then computed as a weighted average of bin centers, yielding the final $\boldsymbol{\lambda}\in\mathbb{R}^{1\times h\times w}$ map. This branch is supervised via the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$. In the Zero Inflation branch (bottom), the same features are fed into a $\boldsymbol{\pi}$-head with a sigmoid activation to estimate the structural zero probability map $\boldsymbol{\pi}\in\mathbb{R}^{1\times h\times w}$. The predicted density map $\boldsymbol{Y}^{*}$ is defined as the expected value of the zero-inflated Poisson distribution: $\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, where $\otimes$ denotes elementwise multiplication. The overall model is trained using the ZIP-based negative log-likelihood loss $\mathcal{L}_{\mathrm{NLL}}$ and an auxiliary count loss $\mathcal{L}_{\mathrm{count}}$ to ensure accurate total count prediction.
Our proposed framework, EBC-ZIP, is an improved extension of the Enhanced Block Classification (EBC) framework introduced in our prior work [15]. In this work, we propose to incorporate Zero-Inflated Poisson (ZIP) regression to address the extreme sparsity of ground-truth density maps. We hypothesize that the zeros in the ground-truth can be categorized into two types: structural zeros, which correspond to semantically irrelevant regions, and sampling zeros, which arise from point-based annotation uncertainty. ZIP regression provides a principled way to model this dual-source sparsity and aligns better with the discrete, non-negative nature of count data. For completeness, we first provide a brief overview of the original EBC framework, followed by a detailed description of ZIP regression. We then present our full EBC-ZIP architecture, which combines the benefits of blockwise classification and probabilistic count modeling. An overview of the proposed framework EBC-ZIP is shown in Fig. 3.
III-A Recap: EBC
Training crowd counting models directly via regression suffers from several limitations. For instance, regression models must learn to accurately predict rare but high-impact values, which can lead to unstable gradient updates – especially in the presence of outliers within a batch. Moreover, regression outputs are typically unbounded (e.g., from a linear layer), which can result in negative or extremely large predictions during early training, causing gradient explosion and slow convergence. The Enhanced Block Classification (EBC) framework addresses these issues by reformulating crowd counting as a blockwise classification problem.
In EBC, blockwise count values are quantized into integer-valued bins $\{\mathcal{B}_{k}\}_{k=1}^{n}$, with larger counts grouped into a shared bin to reduce label sparsity and mitigate the effect of annotation errors. For example, in the case of block size $r=8$, EBC utilizes 00{0}, 11{1}1, 22{2}2, 33{3}3, and $[4,\infty)$ as bins.
Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a backbone network $\mathcal{F}$ (e.g., ResNet [31], VGG [32]) is used to extract feature maps:
where $C$ is the number of channels, and $h=H//r$, $w=W//r$ correspond to the blockwise spatial resolution given block size $r$.
A $1\times 1$ convolutional layer $\mathcal{H}$ followed by softmax is applied to produce the predicted blockwise classification probability map:
$\boldsymbol{P}^{*}=\mathrm{Softmax}\left(\mathcal{H}(\boldsymbol{f})\right)\in\mathbb{R}^{n\times h\times w}.$
Here, the vector $\boldsymbol{P}^{*}_{:,i,j}\in\mathbb{R}^{n}$ denotes the categorical distribution over the $n$ predefined bins for block $(i,j)$.
The final predicted blockwise density map $\boldsymbol{Y}^{*}$ is obtained by taking the expectation over bin centers ${b_{k}}_{k=1}^{n}$:
$\boldsymbol{Y}^{*}_{i,j}=\sum_{k=1}^{n}\boldsymbol{P}^{*}_{k,i,j}\cdot b_{k}.$ (2)
While EBC alleviates the challenges of regression through classification, it still does not fully address the extreme sparsity of ground-truth density maps, where most blocks have zero count (see Fig. 1). Also, its DMCount-based loss function contradicts the discrete nature of blockwise counts. To handle these issues, we introduce zero-inflated Poisson (ZIP) regression, which forms the basis of our proposed extension, EBC-ZIP.
@@ III-A Recap: EBC
Figure 4: Illustration of structural and sampling zeros (best viewed in color). The image (synthesized for illustration purposes) is overlaid with a grid where each cell represents a spatial block. The red dot marks a ground-truth head annotation, and the block containing it is labeled with the red number “1”. Yellow zeros, which indicate sampling zeros, are assigned to blocks that correspond to the head-center region but receive zero count due to the point-based annotation protocol and randomness. Green zeros denote structural zeros, corresponding to background regions, non-head body parts or outer regions of the head that are not associated with any annotations.
@@
Figure 5: Performance of the VGG19-based structure on ShanghaiTech B under varying values of $\omega$ in the loss function (15). Results indicate that $\omega=0.75$ yields the best MAE (6.34) and RMSE (9.79), while $\omega=1.00$ gives the lowest NAE (5.00%).
III-B Zero-Inflated Poisson Regression
Poisson regression is a classical method for modeling count data, where the observed variable $u\in\mathbb{N}_{+}$ is assumed to follow a Poisson distribution with rate parameter $\lambda>0$. The probability mass function (p.m.f.) is defined as:
$P(u\mid\lambda)=\frac{\lambda^{u}e^{-\lambda}}{u!},\quad u\in\{0,1,2,\cdots\}$ (3)
The Poisson distribution is widely used to model the number of discrete events occurring within a fixed interval of time or space, such as the the number of cars passing through a checkpoint per minute, or the number of phone calls received by a call center in an hour. In the context of crowd counting, it is natural to model the number of people appearing in a spatial region (e.g., an image block) as a Poisson random variable, where $\lambda$ denotes the expected count. To reasonably assume independence between blocks, the block size must be sufficiently small, since person-to-person interactions tend to be local. In contrast, using large regions (such as in early works like [29]) may violate this assumption, as spatial correlations between regions may increase in such a scenario. Consequently, it is inappropriate to represent the overall likelihood as the product of probability mass functions.
However, standard Poisson distribution underestimates the probability of observing zeros when the data is zero-heavy. To explicitly handle this, we adopt the Zero-Inflated Poisson (ZIP) model, which which introduces a latent binary process to account for structural sparsity. The ZIP distribution is a mixture of a point mass at zero and a standard Poisson distribution:
$P(v\mid\pi,\lambda)=\begin{cases}\pi+(1-\pi)e^{-\lambda},\;&v=0\\ (1-\pi)\frac{\lambda^{v}e^{-\lambda}}{v!},\;&v>0\end{cases}$ (4)
Here, $\pi\in[0,1]$ represents the probability of a structural zero, and $\lambda$ is the rate of the Poisson component responsible for modeling counts. The probability of a sampling zero is given by $(1-\pi)\exp({-\lambda})$.
This decomposition aligns well with the inherent structure of crowd counting data. Most zero-valued blocks in ground-truth density maps correspond to semantically irrelevant regions such as background (e.g., sky, buildings), non-head body parts (e.g., torsos, limbs), or peripheral head regions, which are never annotated. We refer to these as structural zeros, and explicitly modeling them helps the model suppress spurious count predictions in these regions. In contrast, some zero-valued blocks occur in semantically valid regions (e.g., the head-center area of a person) but do not receive annotation due to the limitations of point-based labeling: Since each person is annotated using a single point (typically the head center), and a head may cover multiple blocks, only one of these blocks receives a non-zero label, while others remain zero. Additionally, annotator subjectivity can cause annotation points to fall in different blocks across annotators. These blocks are not truly empty but are instead artifacts of annotation sparsity and subjectivity, and are therefore better modeled as sampling zeros, arising from the Poisson component. Fig. 4 illustrates these two types of zeros.
The expectation of $v\sim\mathrm{ZIP}(\pi,\lambda)$ can be derived as:
$\mathbb{E}[v\mid\pi,\lambda]$ $=\sum_{k=0}^{\infty}k\cdot P(v=k\mid\pi,\lambda)$
$=\sum_{k=1}^{\infty}k\cdot P(v=k\mid\pi,\lambda)$
$=(1-\pi)\sum_{k=0}^{\infty}k\cdot\frac{\lambda^{k}e^{-\lambda}}{k!}$
$=(1-\pi)\mathbb{E}[u\mid\lambda]=(1-\pi)\lambda,$ (5)
where $u\sim\mathrm{Poisson}(\lambda)$ and $\mathbb{E}(u\mid\lambda)=\lambda$. This expectation confirms that the predicted count is suppressed proportionally to the probability of structural zeros, yielding a more faithful modeling of blockwise density in sparse crowd scenes.
III-C Extending EBC to EBC-ZIP
@@
Similar to EBC, our proposed EBC-ZIP framework is backbone-agnostic, allowing it to be integrated with various feature extractors. Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, we first compute the feature map $\boldsymbol{f}$ using a shared backbone, as described in Eq. (1). We then adopt the EBC framework to predict Poisson rates, due to its training stability and robustness to label errors. Specifically, the feature map $\boldsymbol{f}$ is passed through a $\boldsymbol{\lambda}$-head $\mathcal{H}_{\boldsymbol{\lambda}}$ (implemented as a $1\times 1$ convolution), followed by a softmax activation to generate a probability distribution over the $n{-}1$ positive count bins:
@@
$\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}=\mathrm{Softmax}\left(\mathcal{H}_{\boldsymbol{\lambda}}(\boldsymbol{f})\right)\in\mathbb{R}^{(n-1)\times h\times w}$ (6)
@@
where $h=H//r$, $w=W//r$ denote the height and width of the blockwise output map. Note that we exclude the $\{0\}$ bin here, since the Poisson distribution assumes a strictly positive rate $\lambda>0$. We then compute the estimated Poisson rate map $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{1\times h\times w}$ as the expected value of the bin centers, analogous to Eq. (2):
$\boldsymbol{\lambda}_{i,j}=\sum_{k=1}^{n-1}{\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}}_{k,i,j}\cdot{b_{+}}_{k},$ (7)
where ${b_{+}}_{k}$ for $k=1,\cdots,n-1$ denotes the center of each positive bin.

### v2

# ZIP: Scalable Crowd Counting via Zero‑Inflated Poisson Modeling

@@ Abstract
Most crowd counting methods directly regress blockwise density maps using Mean Squared Error (MSE) losses. This practice has two key limitations: 1) it fails to account for the extreme spatial sparsity of annotations—over 95% of $8\times 8$ blocks are empty across standard benchmarks, so supervision signals in informative regions are diluted by the predominant zeros; 2) MSE corresponds to a Gaussian error model that poorly matches discrete, non‑negative count data. To address these issues, we introduce ZIP, a scalable crowd counting framework that models blockwise counts with a Zero‑Inflated Poisson likelihood: a zero‑inflation term learns the probability a block is structurally empty (handling excess zeros), while the Poisson component captures expected counts when people are present (respecting discreteness). We provide a generalization analysis showing a tighter risk bound for ZIP than MSE‑based losses and DMCount provided that the training resolution is moderately large. To assess the scalability of ZIP, we instantiate it on backbones spanning over 100×100$\times$100 × in parameters/compute. Experiments on ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd demonstrate that ZIP consistently surpasses state‑of‑the‑art methods across all model scales.
Introduction
@@
Crowd counting aims to estimate the number of people present in an image or a video frame. It underpins a broad range of real-world applications, including public safety and crowd management (Valencia et al. 2021; Saleh, Suandi, and Ibrahim 2015), urban planning (Waclawek et al. 2024), and intelligent transportation systems (McCarthy et al. 2025).
Most modern crowd counting methods (Ma et al. 2019; Cheng et al. 2022; Han et al. 2023; Ranasinghe et al. 2024) construct blockwise training targets by first convolving point annotations with Gaussian kernels to create a pixel‑level density map, and then summing density values within each $B\times B$ block (e.g., $8\times 8$) to obtain blockwise density maps. Models are trained to regress these blockwise targets (usually with Mean Squared Error (MSE)) in lieu of raw pixel supervision to mitigate extreme pixel‑level sparsity.
In these methods, however, the aggregated blockwise density maps remain highly sparse: across ShanghaiTech A & B (Zhang et al. 2016), UCF‑QNRF (Idrees et al. 2018), and NWPU‑Crowd (Wang et al. 2020c), over 95% of $8\times 8$ blocks contain zero people (empirical measurement). This zero dominance skews squared‑error losses, dilutes supervision from populated regions, and biases models toward under‑counting. These methods also suffer label noise introduced by Gaussian smoothing. Point annotations are spatially uncertain: a human click marks an approximate head center, and nearby pixels could be equally valid. Spreading each point with a local kernel is intended to make supervision tolerant to such subjective variation. However, since the true head extents are unknown, mismatched kernel bandwidths yield over‑ or under‑counts after block aggregation and thus inject noise into learning targets.
(a) Input Image.
(b) GT Density Map.
(c) Structural Zero Map.
(d) Predicted Density Map.
Figure 1: Visualization of ZIP on a ShanghaiTech B test image. Color scale (all panels): red = higher value, blue = lower value (per-panel normalization). In Fig. 1(c), red regions have high probability of being structurally empty (background, torso, peripheral head parts), blue indicates low structural-zero probability (candidate head centers). In Fig. 1(d), red regions correspond to high expected local crowd density near head centers; blue denotes near-zero density.
Prior work has tried to mitigate sparsity by first segmenting foreground crowd regions and then counting only within those areas (Rong and Li 2021; Modolo et al. 2021; Guo et al. 2024). Such a practice introduces additional multi-task loss balancing complexity. Also, since crowd counting datasets are not annotated with segmentation masks, these approaches usually derive pseudo masks from Gaussian-smoothed density maps, inheriting the same head-size uncertainty and label noise discussed above.
@@
Figure 2: Illustration of structural and sampling zeros (best viewed in color). The image (synthesized for illustration purposes) is overlaid with a grid where each cell represents a spatial block. The red dot marks a ground-truth head annotation, and the block containing it is labeled with the red number “1”. Yellow zeros, which indicate sampling zeros, are assigned to blocks that correspond to the head-center region but receive zero count due to the point-based annotation protocol. Green zeros denote structural zeros, corresponding to background regions, non-head body parts or outer regions of the head that are not associated with any annotations.
Additionally, MSE-based loss functions present a modeling mismatch. Blockwise targets represent discrete, non‑negative, and zero-heavy count data, whereas MSE implicitly assumes Gaussian residuals on continuous values. A Poisson model captures both the discreteness and non-negativity of counts, yet a standard Poisson struggles with the overwhelming presence of zeros. These gaps motivate a formulation that explicitly separates structural emptiness from stochastic counts, leading to our Zero-Inflated Poisson (ZIP) framework.
Our ZIP framework bypasses Gaussian smoothing and model per-block counts $y\in\mathbb{N}$ via a zero‑inflated Poisson:
$P(y=k|\pi,\lambda)=\begin{cases}\pi+(1-\pi)e^{-\lambda},\quad&k=0\\ (1-\pi)\frac{e^{-\lambda}\lambda^{k}}{k!},\quad&k>0\end{cases}$ (1)
with $\pi$, $\lambda$ predicted per block. In this way, we hypothesize that zeros in blockwise density maps arise from two distinct mechanisms (See Fig. 2 for illustration):
  • Structural zeros (modeled by $\pi$), which are blocks that are deterministically zero due to their semantic irrelevance. These include background (sky, buildings, etc.), body parts (limbs, chest, etc.) and peripheral head regions that do not correspond to head centers, and they comprise the majority of zero blocks.
  • Sampling zeros (modeled by $(1-\pi)e^{-\lambda}$): each head is annotated by a single point that we assign to exactly one supervision block. Neighboring blocks that also correspond to the head center therefore receive zero count. Modeling these annotation effects in the Poisson component absorbs the positional ambiguity that Gaussian smoothing was meant to solve, removing kernel bandwidth tuning.
Because structural and sampling zeros are not annotated separately, our ZIP model learns to distinguish these two types of zeros implicitly by optimizing the negative log-likelihood of the ZIP distribution, with $\lambda$ additionally supervised by a cross entropy loss to stabilize training. Empirically, learned zero‑inflation correlates with non-head-center areas. As illustrated in Fig. 1, the structural-zero map in 1(c) predicts nearly all non-head pixels as zeros, while preserving only the compact blue blobs centered on heads. Compared with segmentation-based methods, our ZIP-based method offers several notable advantages: 1) it does not rely on segmentation masks, thereby avoiding noise introduced by Gaussian smoothing; 2) it captures structural sparsity without assuming the presence of clear object boundaries, making it better suited for highly occluded and densely crowded scenes; 3) it eliminates the need to balance counting and segmentation, simplifying the training objective.
Figure 3: Scalability of the proposed ZIP framework. The $x$-axis represents the number of model parameters (in millions), and the $y$-axis indicates the Mean Absolute Error (MAE; lower is better) on the validation set of the ShanghaiTech B dataset (Zhang et al. 2016). Each circle represents a specific model, where the radius corresponds to the number of floating point operations (FLOPs) required to process a single $1920\times 1080$ image during inference. ZIP models (in blue) exhibit a favorable trade-off between model size and accuracy, demonstrating better scalability compared to other lightweight (yellow) and heavyweight (green) models.
Through theoretical risk bound analysis (See Theorem 1) and comprehensive empirical evaluation, we show that ZIP consistently outperforms Gaussian-smoothed MSE-based losses and DMCount (Wang et al. 2020a). To further assess the scalability of ZIP, we evaluate its performance with backbones of varying computational complexities, including MobileNet, MobileCLIP , and CLIP-ConvNeXt. As illustrated in Fig. 3, our ZIP framework consistently outperforms existing models in both accuracy (measured by MAE) and computational efficiency (measured by FLOPs) across different parameter scales, demonstrating its superior scalability. To summarize, our contributions are as follows:
  • We propose a novel crowd counting framework ZIP based on Zero-Inflated Poisson (ZIP) regression, which explicitly models the extreme sparsity of ground-truth count distributions by disentangling structural zeros from sampling uncertainty.
  • We conduct a theoretical analysis and experiments on four crowd counting benchmarks to validate the effectiveness of our ZIP framework. Results show that our base model, ZIP-Base, consistently outperforms existing state-of-the-art methods across four different datasets.
  • We present a systematic study of framework-level scalability in crowd counting. Our ZIP framework generalizes well across a wide range of model sizes and architectures, from lightweight convolutional neural networks (CNNs) to vision-language models, achieving state-of-the-art performance under varying computational constraints.
Related Work
To address the extreme sparsity of ground-truth density maps, many existing methods (Zhang et al. 2016; Liu, Salzmann, and Fua 2019; Ma et al. 2019; Wang et al. 2020b, 2021; Han et al. 2023; Guo et al. 2024) adopt Gaussian smoothing. These approaches typically preprocess the ground-truth density map by convolving it with Gaussian kernels, either of fixed or adaptive size. In principle, the kernel size should reflect the actual head size in the image; however, such information is typically unavailable in most crowd counting datasets. As a result, Gaussian smoothing inevitably introduces noise into the ground-truth density maps, which can degrade the model’s generalization ability (Wang et al. 2020a). There have also been some efforts (Rong and Li 2021; Modolo et al. 2021; Guo et al. 2024) to address the sparsity by training a segmentation model which aims to separate regions containing people from the background. However, most crowd counting datasets do not provide segmentation masks, they have to rely on Gaussian-smoothed ground-truth density maps to generate pseudo ground-truth segmentation masks. This again introduces noise into the pseudo segmentation masks. Guo et al. exploited the box annotations provided by NWPU-Crowd (Wang et al. 2020c) to handle this problem, but compared with point annotations, box annotations are much more expensive to acquire. Besides, these methods introduce an extra task to counting, and how to balance the loss terms in such a multi-task setting poses a new challenge.
Figure 4: Overview of the proposed ZIP framework. Given an input image $\boldsymbol{X}$ of spatial size $H\times W$, a backbone network extracts feature maps, which are shared by two parallel branches. In the Poisson branch (top), the features are passed to a $\boldsymbol{\lambda}$-head with a softmax activation, producing the blockwise count distribution $\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}$ over $(n-1)$ positive bins. The Poisson rate per block is then computed as a weighted average of bin centers, yielding the final $\boldsymbol{\lambda}$ map. In the Zero Inflation branch (bottom), the same features are fed into a $\boldsymbol{\pi}$-head with a sigmoid activation to estimate the structural zero probability map $\boldsymbol{\pi}$ of spatial size $h\times w$. The predicted density map $\boldsymbol{Y}^{*}$ is defined as the expected value of the zero-inflated Poisson distribution: $\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, where $\otimes$ denotes elementwise multiplication.
On the other hand, Most density-based crowd counting methods typically frame the task as a regression problem, where the model is trained to minimize the blockwise mean squared error (MSE) between the predicted and ground-truth blockwise density maps (Zhang et al. 2016; Liu, Salzmann, and Fua 2019; Liu et al. 2019; Ma et al. 2019; Wang et al. 2020b; Han et al. 2023; Guo et al. 2024). This loss function implicitly assumes that each blockwise count follows a Gaussian distribution centered at the predicted value. This formulation fails to capture the discreteness and non-negativity of count data. DMCount (Wang et al. 2020a) addresses a different issue by reformulating density estimation as an Earth Mover’s Distance (EMD) problem. By minimizing the Wasserstein distance between predicted and ground-truth density maps, DMCount explicitly models the global transportation cost needed to match the predicted mass to the annotated points. Similar to MSE-based methods, it still operates entirely in a continuous density space, treating the density maps as divisible mass rather than integer-valued counts. Also, the exact computation of EMD scales poorly with the number of bins ($\mathcal{O}(n^{3})$).
Method
In this work, we propose to incorporate Zero-Inflated Poisson (ZIP) regression to address the extreme sparsity of ground-truth density maps. Fig. 4 illustrates an overview of the proposed ZIP framework.
Zero-Inflated Poisson Regression
Since training crowd counting models directly via regression can suffer from unstable gradient updated, we follow the Enhanced Block Classification (EBC) framework (Ma, Sanchez, and Guha 2024) to first quantize blockwise counts into integer-valued bins $\{\mathcal{B}_{k}\}_{k=1}^{n}$ with bin centers $\{b_{k}\}_{k=1}^{n}$.
@@
Figure 5: Performance of the VGG19-based structure on ShanghaiTech B under varying values of $\omega$ in the loss Eq. (11). Results indicate that $\omega=0.75$ yields the best MAE (6.34) and RMSE (9.79), while $\omega=1.00$ gives the lowest NAE (5.00%).
Given an input image $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, we first compute the feature map $\boldsymbol{f}$ using a shared backbone:
where $h=H//r$, $w=W//r$ denote the spatial size of the output.
@@
Then, the feature map $\boldsymbol{f}$ is passed through a $\boldsymbol{\lambda}$-head $\mathcal{H}_{\boldsymbol{\lambda}}$ (implemented as a $1\times 1$ convolution), followed by a softmax activation to generate a probability distribution over the $n{-}1$ positive count bins:
@@
$\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}=\mathrm{Softmax}\left(\mathcal{H}_{\boldsymbol{\lambda}}(\boldsymbol{f})\right)\in\mathbb{R}^{(n-1)\times h\times w}.$ (3)
@@
Note that we exclude the $\{0\}$ bin here, since the Poisson distribution assumes a strictly positive rate $\lambda>0$. We then compute the estimated Poisson rate map $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{1\times h\times w}$ as the expected value of the bin centers:
$\boldsymbol{\lambda}_{i,j}=\sum_{k=2}^{n}{\boldsymbol{P}^{*}_{\boldsymbol{\lambda}}}_{k,i,j}\cdot{b}_{k},$ (4)
where ${b}_{k}$ for $k=2,\cdots,n$ denotes the center of each positive bin.